OpenStudy (anonymous):

Solve this inequality. -3x - 2 > 7 x < 5 -3 x > 5 -3 x < -3 x > -3

OpenStudy (ajprincess):

solving an inequality means to hav the variable on one side of the inequality and the numbers on the other side. here x is the variable. to get it alone we need to remove all the numbers associated with that. first let us remove the 2 that is being subtracted from x. to do that add 2. like this -3x-2+2>7+2 remember when u make alteration to one side of the inequality you need to do the same to the other side.

OpenStudy (anonymous):

a \[x \le \frac{ 5 }{ 3 }\] b \[x\ge \frac{ 5 }{-3}\] c\[x \le -3\] d\[x \ge -3\]

OpenStudy (ajprincess):

-3x-2+2>7+2 will result in -3x>9 now u need to remove -3 that has been multiplied to x. to do that divide by -3. remember when u divide by a negative number the inequality sign flips.

OpenStudy (anonymous):

so i looked at the answers and when simplified i get\[x \le -3\] would that be right?

OpenStudy (ajprincess):

yup right:)